# High-order Fusion Graph Contrastive Learning for Recommendation

### v2

@@ 4.2. Overall Performance Comparisons
  • All GCL-based methods such as SimGCL, RecDCL, and BIGCF demonstrate clear superiority over traditional methods, largely due to the self-supervised signals derived from CL. Within this group, HFGCL achieves the best performance. This is mainly attributed to the high quality contrastive views and the ability to fuse different self-supervised signals with CL loss. Additionally, while RecDCL achieves sub-optimal performance $w.r.t.$ Recall@20 on the Amazon-book dataset, the use of the 2048 embedding size may compromise the fairness of comparisons with other methods. In contrast, HFGCL achieves superior recommendation performance by leveraging high-order contrastive for CL, without the need for any data augmentation.
@@ 4.3. Method Variants and Ablation Study
From the results shown in Table 5, we draw the following conclusions. Firstly, HFGCL outperforms LightGCN across all performance metrics. For variant $\mathrm{HFGCL}_{b}$, which utilizes user-item contrastive views, the performance not only matches that of SGL but also significantly surpasses LightGCN. For variant $\mathrm{HFGCL}_{h}$, which eliminates low-order information, the CL loss effectively gets self-supervised signals from high-order information. For variant $\mathrm{HFGCL}_{s}$, with the addition self-sueprvies signals, there is a discernible enhancement in the method’s recommendation quality. These variants adequately address the question raised in Section 2.3 regarding the necessity of data augmentations, demonstrating that it is not essential for creating contrastive views.
Furthermore, for the fusion CL loss, we conduct a series of ablation studies to validate the effectiveness. For $\mathrm{HFGCL_{w/o\,high}}$, the results show that low-order information hinders similarity between positive pairs. For $\mathrm{HFGCL_{w/o\,GCN}}$, $\mathrm{HFGCL_{w/o\,view}}$ and $\mathrm{HFGCL_{w/o\,CL}}$, we remove the widely used GCN encoder, use the original CL loss and remove the CL loss, respectively, and the experimental results demonstrate the power of the fusion CL loss. Finally, compared to HFGCL, all ablations result in significant performance declines, which clearly demonstrate the superiority of the fusion CL loss.
@@ 4.4. Method Efficiency Study
To verify the efficiency of HFGCL, we compare the total training cost of the state-of-the-art methods. As shown in Fig. 5, our HFGCL has the shortest training cost on three datasets. This efficiency is primarily due to HFGCL quickly fuse self-supervised signals over the contrastive views, which enables it to converge exceedingly quickly. For SGL and SimGCL, both data augmentation necessitate graph convolution, a process that is notably time-consuming. For NCL and CGCL, though they also do not require data augmentations, it is again extremely time-consuming to compute all users/items from the user-item interaction as negative samples. Moreover, for most of the GCL-based methods (using data augmentations), the overall convergence speed is significantly impacted because the quality of the generated embedding information is perturbed to varying degrees. It is noteworthy that despite the absence of training cost data for RecDCL, it ranks as the most time-consuming method in the comparison, primarily because of its embedding size of 2048.
@@ 4.5. Method Sparsity Study
To show the ability of HFGCL in addressing the data sparsity issue, we categorize the users of the dataset into three groups—sparse, common, and popular—following methodologies from previous research. The experimental results are shown in Fig. 6. HFGCL maintains superior performance across different sparsity degrees, and performs exceptionally well within the sparsest user category. Additionally, on Amazon-book dataset, we observe a general decrease in performance in the normal category, possibly attributed to high levels of noise in this user segment. Even so, HFGCL maintains superior recommendation performance, demonstrating its robustness and resistance to sparsity.
@@ 4.6. Method Hyperparameter Study
In this section, we provide the sensitivity of HFGCL to various hyperparameters in Fig. 7 and Table 6. As depicted in Fig. 7, the method’s performance is significantly affected by the value of $\tau$, which is slightly higher compared to the previous method. This increase is primarily due to the elevated number of negative samples, necessitating a larger $\tau$ to ensure optimal acquisition of mutual information. However, HFGCL’s performance is barely affected by $\lambda_{1}$. We attribute this to our CL loss’s robust adaptation to the high-order contrastive views, ensuring the method performs optimally regardless of the $\lambda_{1}$ value. This efficacy underscores the strength of HFGCL’s CL loss. The results presented in Table 6 confirm the efficacy of high-order information, demonstrating that HFGCL significantly enhances method performance when $l\geq h$. This also confirms that the low-order information decreases the similarity between positive pairs, leading to suboptimal recommendation.
@@ 5. RELATED WORK
Furthermore, GNNs have been extensively explored in various recommendation scenarios (Wang et al., 2022b; Sharma et al., 2024). In sequential recommendation, many methods, $e.g.,$ SURGE (Chang et al., 2021) and GCE-GNN (Wang et al., 2020), leverage GNN to aggregate items within each sequence via interaction graphs, thereby enhancing the quality of item encoding. In social recommendation, numerous methods, $e.g.,$ DVGRL (Zhang et al., 2024c), GraphRec (Fan et al., 2019), and ESRF (Yu et al., 2022a), combine interaction graphs with social relationship graphs to more effectively incorporate social relationships. However, despite the effectiveness, their lack of self-supervised signals hinders their ability to accurately capture users’ preferences.
CL-based Recommendation. Recent developments in recommender systems research have predominantly emphasized contrastive learning (CL) (Yu et al., 2024; Liu et al., 2023b). The self-supervised signals generated by CL effectively address data sparsity issues. CL-based methods like NCL (Lin et al., 2022), SimGCL (Yu et al., 2022b), VGCL(Yang et al., 2023b), and BIGCF (Zhang et al., 2024b) typically employ data augmentation to create contrastive views, by facilitating CL. For example, SimGCL (Yu et al., 2022b) incorporates noise during graph convolution to generate multiple contrastive views. In contrast to traditional GCL-based methods, NCL (Lin et al., 2022) and CGCL (He et al., 2023) do not rely on data augmentation. For instance, NCL uses cross-layer and clustering methods to generate contrastive views and treats all user and item embeddings as negative samples for training. While these method have some benefits, they significantly increase the training cost. Our HFGCL, however, achieves optimal training efficiency and recommendation performance by considering high-order information as contrastive views and proposing a fusion CL loss.

### v4

@@ 5.2. Overall Performance Comparisons
  • All CL-based methods such as SimGCL, RecDCL, and BIGCF demonstrate clear superiority over traditional methods, largely due to the self-supervised signals derived from CL. Within this group, HFGCL achieves the best performance. This is mainly attributed to the high quality contrastive views and the ability to fuse different self-supervised signals with CL loss. Additionally, while RecDCL achieves sub-optimal performance $w.r.t.$ Recall@20 on the Amazon-book dataset, the use of the 2048 embedding size may compromise the fairness of comparisons with other methods. In contrast, HFGCL achieves superior recommendation performance by leveraging high-order contrastive for CL, without the need for any data augmentation.
@@ 5.3. Method Variants and Ablation Study
From the results shown in Table 7, we draw the following conclusions. Firstly, HFGCL outperforms LightGCN across all performance metrics. For variant $\mathrm{HFGCL}_{b}$, which utilizes user-item contrastive views, the performance not only matches that of SGL-ED but also significantly surpasses LightGCN. For variant $\mathrm{HFGCL}_{h}$, which eliminates low-order information, the CL loss effectively gets self-supervised signals from high-order information. For variant $\mathrm{HFGCL}_{s}$, with the addition self-sueprvies signals, there is a discernible enhancement in the method’s recommendation quality. These variants adequately address the question raised in Section 3.1 regarding the necessity of data augmentations, demonstrating that it is not essential for creating contrastive views.
Furthermore, for the fusion CL loss, we conduct a series of ablation studies to validate the effectiveness. For $\mathrm{HFGCL_{w/o\,high}}$, the results show that low-order information hinders similarity between positive pairs. For $\mathrm{HFGCL_{w/o\,GCN}}$, $\mathrm{HFGCL_{w/o\,view}}$ and $\mathrm{HFGCL_{w/o\,CL}}$, we remove the widely used GCN encoder, use the original CL loss and remove the CL loss, respectively, and the experimental results demonstrate the power of the fusion CL loss.
@@ 5.4. Method Efficiency Study
To verify the efficiency of HFGCL, we compare the total training cost of the state-of-the-art methods. As shown in Fig. 4, our HFGCL has the shortest training cost on three datasets. This efficiency is primarily due to HFGCL quickly fuse self-supervised signals over the contrastive views, which enables it to converge exceedingly quickly. For SGL-ED and SimGCL, both data augmentation necessitate graph convolution, a process that is notably time-consuming. For NCL and CGCL, though they also do not require data augmentations, it is again extremely time-consuming to compute all users/items from the user-item interaction as negative samples. Moreover, for most of the GCL-based methods (using data augmentations), the overall convergence speed is significantly impacted because the quality of the generated embedding information is perturbed to varying degrees. It is noteworthy that despite the absence of training cost data for RecDCL, it ranks as the most time-consuming method in the comparison, primarily because of its embedding size of 2048.
@@ 5.5. Method Sparsity Study
To show the ability of HFGCL in addressing the data sparsity issue, we categorize the users of the dataset into three groups—sparse, common, and popular—following methodologies from previous research. The experimental results are shown in Fig. 5. In the sparsity study, we focus on the performance of the method in the sparse user group. Compared to existing SOTA methods, SimGCL and BIGCF, HFGCL shows significant performance improvements across all three datasets. Additionally, HFGCL maintains excellent performance across different levels of sparsity. On the Amazon-book dataset, we observe a general decline in performance for the regular user category, possibly due to higher noise levels in this group. Nevertheless, HFGCL still demonstrates strong recommendation performance, proving its robustness and resistance to sparsity.
@@ 5.6. Method Hyperparameter Study
In this section, we provide the sensitivity of HFGCL to various hyperparameters in Fig. 6 and Table 8. As depicted in Fig. 6, the method’s performance is significantly affected by the value of $\tau$, which is slightly higher compared to the previous method. This increase is primarily due to the elevated number of negative samples, necessitating a larger $\tau$ to ensure optimal acquisition of mutual information. However, HFGCL’s performance is barely affected by $\lambda_{1}$. We attribute this to our CL loss’s robust adaptation to the high-order contrastive views, ensuring the method performs optimally regardless of the $\lambda_{1}$ value. This efficacy underscores the strength of HFGCL’s CL loss. Table 8 shows the performance results when the high-order initialization layer $h$ of HFGCL is set to 1, 2 and 3. The results confirm the effectiveness of high-order information, indicating that when $h\leq l\leq L$, HFGCL significantly improves performance. It also shows that stacking low-order information causes nodes to contain excessive self-information, reducing the similarity between positive pairs and resulting in suboptimal recommendations.
@@ 6. RELATED WORK
Furthermore, GNNs have been extensively explored in various recommendation scenarios (Wang et al., 2022b; Sharma et al., 2024). In sequential recommendation, many methods, $e.g.,$ GCE-GNN (Wang et al., 2020) leverage GNN to aggregate items within each sequence via interaction graphs, thereby enhancing the quality of item encoding. In social recommendation, numerous methods, $e.g.,$ DVGRL (Zhang et al., 2024c) and GraphRec (Fan et al., 2019) combine interaction graphs with social relationship graphs to more effectively incorporate social relationships. However, despite the effectiveness, their lack of self-supervised signals hinders their ability to accurately capture users’ preferences.
CL-based Recommendation. Recent developments in recommender systems research have predominantly emphasized contrastive learning (CL) (Yu et al., 2024). The self-supervised signals generated by CL effectively address data sparsity issues. CL-based methods like NCL (Lin et al., 2022), SimGCL (Yu et al., 2022), VGCL(Yang et al., 2023b), and BIGCF (Zhang et al., 2024b) typically employ data augmentation to create contrastive views, by facilitating CL. For example, SimGCL (Yu et al., 2022) incorporates noise during graph convolution to generate multiple contrastive views. In contrast to traditional GCL-based methods, NCL (Lin et al., 2022) and CGCL (He et al., 2023) do not rely on data augmentation. For instance, NCL uses cross-layer and clustering methods to generate contrastive views and treats all user and item embeddings as negative samples for training. While these method have some benefits, they significantly increase the training cost. Our HFGCL, however, achieves optimal training efficiency and recommendation performance by considering high-order information as contrastive views and proposing a fusion CL loss.